The Greater Pittsburgh Area Chamber of Commerce wants to estimate the mean time workers who are employed in the downtown area spend getting to work. A sample of 15 workers reveals the following number of minutes spent traveling. 29 38 38 33 38 21 45 34 40 37 37 42 30 29 35 a. Develop a 95 percent confidence interval for the population mean. Interpret the result. 


1
Expert's answer
2021-01-12T14:55:33-0500

Sample mean

"μ= \frac{Sum \;of\; all \; the \; observations}{Number \; of \; observations} \\

= \frac{29+38+38+33+38+21+45+34+40+37+37+42+30+29+35}{15} \\

= \frac{526}{15} \\

= 35.067"

Standard deviation (SD) is formulated as follows:

"SD = \sqrt{\frac{1}{n-1}\sum(x_i-μ)^2}"


"SD = \sqrt{\frac{507}{15-1}} \\

= \sqrt{36.21} \\

= 6.017"

Now we will develop a 95 confidence interval.

Use the z-value for 95%: z = 1.96

Standard error is given as:

"SE = \frac{z \times SD}{\sqrt{n}} \\

= \frac{1.96 \times 6.017}{\sqrt{15}} \\

= 3.045"

Confidence interval is:

"(μ-SE, μ+SE) \\

(35.067-3.045, 35.067+3.045) \\

(32.022, 38.112)"

The mean time workers who are employed in the downtown area spend getting to work is between 32.022 and 38.112.